# Tera-MIND: Tera-scale mouse brain simulation via spatial mRNA-guided diffusion

### v2

@@
Abstract
@@
Holistic 3D modeling of molecularly defined brain structures is crucial for understanding complex brain functions. Emerging tissue profiling technologies enable the construction of a comprehensive atlas of the mammalian brain with sub-cellular resolution and spatially resolved gene expression data. However, such tera-scale volumetric datasets present significant computational challenges in understanding complex brain functions within their native 3D spatial context. Here, we propose the novel generative approach Tera-MIND, which can simulate Tera-scale Mouse braINs in 3D using a patch-based and boundary-aware Diffusion model. Taking spatial transcriptomic data as the conditional input, we generate virtual mouse brains with comprehensive cellular morphological detail at teravoxel scale. Through the lens of 3D gene-gene self-attention, we identify spatial molecular interactions for key transcriptomic pathways in the murine brain, exemplified by glutamatergic and dopaminergic neuronal systems. Importantly, these in-silico biological findings are consistent and reproducible across three tera-scale virtual mouse brains. Therefore, Tera-MIND showcases a promising path toward efficient and generative simulations of whole organ systems for biomedical research. Project website: https://musikisomorphie.github.io/Tera-MIND.html
1 Introduction
@@
Recently, a network of researchers [47, 49, 21, 35] has reached a groundbreaking milestone by charting the first complete cellular atlas of the adult mouse brain, the most extensively studied animal model in neuroscience. Using advanced spatial transcriptomics methods [3, 25, 5] capable of profiling molecularly defined brain anatomy and morphology at sub-cellular resolution, the collective studies established foundational resources and datasets, providing an unprecedented framework for systematically exploring the molecular-driven spatial organization of the mammalian brain. Consequently, this landmark achievement opens the door for next-generation investigations into intricate brain functions characterized by exceptional spatial complexity. However, these tera-scale volumetric datasets, including spatial molecular data (e.g., mRNA readouts acting as proxies for gene expression) and paired morphological bioimages, pose substantial computational challenges [45] for in-silico modeling of the mouse brain. Integrating spatial gene expression with 3D brain morphology to uncover functional relationships at whole organ scale remains an open challenge.
@@
Meanwhile, generative artificial intelligence (GenAI) has emerged as a transformative tool for automating visual content creation and manipulation [4, 48]. Given a text prompt that describes the desired visual content, GenAI models [14, 15] are able to produce highly realistic images and videos [32, 30] with remarkable efficiency. In biomedical research, the applications of GenAI have gained domain-specific recognition in various areas such as digital histopathology [2], drug screening [43, 20], and molecule design [4]. However, the direct application of existing GenAI methods to whole organ simulation has currently remained infeasible due to technological limitations. Developed upon the framework of Generative Adversarial Nets (GAN) [14], the GigaGAN [19] model was trained on 96-128 A100 GPUs to synthesize $096\times 4,$ 096 × 4 , 096 color images. Using up to 6,144 H100 GPUs, Movie Gen [31], a diffusion-based model [15], achieved video generation results with $1,092\times 1,080\times 256$ spatial-temporal resolution. In comparison, even after image registration and quality control, the mouse brain atlas remains at tera-scale volume, consisting of 100 whole slide images (WSIs; $0.108\mu m/\mathrm{pixel}$) paired with size-matched gene expression arrays, each with a spatial resolution of $73,216\times 105,984$. This immense data scale of WSIs (e.g., $73,216\times 105,984\times 100=0.77\times 10^{12}$ voxels) as well as gene expression arrays significantly amplifies the computational demands associated with GenAI modeling of a virtual mouse brain.
@@
Addressing memory bottlenecks in high-resolution image generation has previously motivated the development of patch-based approaches [22, 23, 46, 28, 12], which enforce boundary consistency [45] between generated neighboring patches. Similar to many standard GenAI models, these approaches can also leverage the interplay between textual and visual modalities and create desired visual content conditioned on text prompts. While this design choice has shown remarkable success in natural image synthesis and artistic content creation, it is suboptimal for biomedical applications – particularly for the accurate 3D molecular-driven reconstruction of mammalian brains. The inherent complexity of brain morphology controlled by spatial molecular interactions demands a generative framework that can capture molecular-to-morphology spatial associations with precision. In response to these challenges, we propose Tera-MIND, a novel GenAI approach designed for the paired data structure of spatial mRNA arrays (gene expression, prompts) and WSIs (brain morphology, bioimages). Specifically, Tera-MIND employs a newly introduced patch-based and boundary-aware diffusion model, which allows the reconstruction of tera-scale mouse brains with high fidelity. Owing to the patch-wise training and inference paradigm, Tera-MIND models in-silico mammalian brain(s) with computational efficiency. As a result, the whole simulation process can be efficiently executed on a single DGX A100 machine, significantly lowering hardware demands while maintaining scalability for tera-scale data processing. This efficiency underscores the potential wide applicability of Tera-MIND for biomedical applications. Our contributions can be summarized as follows:
@@
  • Lastly, the in-silico biological findings obtained by Tera-MIND can be efficiently reproduced across three tera-scale virtual mouse brains. This confirms the robustness and reliability of Tera-MIND for biomedical applications.
Figure 1: Overview of the Tera-MIND architecture and workflows for virtual organ simulation (Main result). a. The 3D reference map of major brain regions (left) alongside their corresponding labels (right) for guiding the navigation of brain structures. b. The 3D visualization of the ground-truth (GT, left) and generated (right) mouse brain, as captured by a sequential stack of DAPI (4’,6-diamidino-2-phenylindole) and PolyT-stained (total mRNA signals) WSIs at the scale of $0.77\times 10^{12}$ voxels. The bottom part of generated and GT results overlay brain region maps, aiding in the navigation of the complex spatial brain architecture. c. The conceptual illustration of patch-based diffusion model training. Here, we train Tera-MIND with noisy DAPI and PolyT bioimage patches so that the model learns to output clean patches guided by paired spatial mRNA array. Apart from the standard reconstruction path (black arrows), we supply the boundary-aware path (gray arrows) to impose boundary consistency between generated neighboring patches. d. The conceptual illustration of patch-wise generating virtual mouse brain using Tera-MIND. At this stage, we only run through the boundary-aware path to seamlessly generate the tera-scale mouse brain.
2 Results
2.1 Model overview of Tera-MIND
To resolve the high-volume computational demand, we develop a novel patch-based and boundary-aware diffusion model Tera-MIND, designed for the scalable and seamless generation of teravoxel mouse brains (See Fig. 1 (a, b)). Considering the spatial linkage of gene expression variability and brain morphology, we propose to modulate a 3D gene-morph (gm) UNet using a tailored 3D gene-gene (gg) block. By feeding the spatial gene expression array (input prompt) to the 3D-gg block, we achieve the sub-cellular control of brain morphology generation through the 3D-gm UNet, please see Fig. 1 (c) for more illustrations. To assess the role of individual genes and their relationships within gene groups, the proposed 3D gene-gene block includes a 3D self-attention layer that learns the gene-gene interaction level within the native 3D spatial context. Following the attention architecture introduced in Diffusion Transformer (DiT) [30], we then inject learned gene-gene attention representations into the 3D gene-morph cross-attention block, eventually modulating brain morphology generation using 3D-gm UNet. In addition to the patch-wise reconstruction path (black arrows in Fig. 1 (c, d)), we further introduce a boundary-aware path (gray arrows in Fig. 1 (c, d), please see also [12]) to output center-cropped brain morphology patches. Specifically, the training process involves randomly cropping paired $n$-plex gene expression arrays and their corresponding morphological bioimages from training mouse brain atlas(es) [47]. Our patch-based approach thus enables efficient training of the Tera-MIND model on 2 $\times$ 40GB A100 GPUs. During inference, Tera-MIND supports the seamless generation of unseen mouse brain(s) with high fidelity, while the hardware requirement remains moderate, e.g., 7 days on a single DGX machine with 8$\times$ 40GB A100 GPUs. For model details we refer interested readers to the Methods section and code repository.
@@ 2.1 Model overview of Tera-MIND
2.2 Tera-MIND enables accurate generation of tera-scale mouse brain(s) by spatial gene expression.
@@
Consistent with the primary analysis described in [49, 47], which was conducted on a comprehensive brain atlas of a P56 female adult mouse, we present main simulation results of Tera-MIND based on the same brain atlas. The holistic 3D comparison between the ground truth (GT) and our generation result is illustrated in Fig. 1 (b) and Fig. 2 (a). Given 3D spatial gene expression data as an input, Tera-MIND achieves the faithful and comprehensive reconstruction of 3D brain morphology, as captured by a sequential stack of DAPI and PolyT-stained WSIs. Specifically, the generated mouse brain has the same teravoxel scale as GT, e.g., $73,216\times 105,984\times 100=0.77$ teravoxels. At the level of individual WSIs (Fig. 2 (b, c)), Tera-MIND exhibits high-quality generation across multiple scales, including cellular, regional, and slice-wise morphological organization. Within distinct brain regions, such as the olfactory areas, isocortex, and cerebellar cortex, Tera-MIND accurately reconstructs complex morphological structures while preserving native cellular distributions. These results highlight the robustness of Tera-MIND for reconstructing 3D brain morphology directly from spatial gene expression profiles.
@@
Complementary to the side-by-side qualitative comparisons, we perform thorough quantitative analyses of the reconstruction quality by Tera-MIND. For a fair and systematic assessment of the biomedical-specific outputs, we report not only commonly used metrics such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM), but also include patch-based morphometric analyses of nuclear size and cell number. These domain-specific metrics are critical for evaluating the biological fidelity of reconstructed brain morphology at the cellular level. Then, we perform quantitative evaluation on Tera-MIND in comparison to state-of-the-art (SOTA) methods CoCoGAN [22], InfinityGAN [23], IST-editing [42], SinFusion [28], Patch-DM [12], etc. Overall, Tera-MIND achieved superior performance in terms of better PSNR and SSIM scores. Furthermore, experimental results reported in Fig. 2 (e) demonstrated marginal morphometric discrepancies between GT cellular patches and those generated by Tera-MIND, which are consistent with the results of stratified patch examples shown in Fig. 2 (e). Using spatial gene expression as the input prompt, these analyses confirm that Tera-MIND reliably reconstructs brain morphology across multiple scales, from cellular and regional structures to whole-brain organization. This consistent performance across scales highlights the efficacy of our approach for accurately capturing the complex spatial patterns of 3D brain morphology.
Figure 2: The visual and quantitative comparison of mouse brain generation (Main result). a. The 3D map of major brain regions (left), GT (middle), and the generated mouse brain (right) of the atlas used in the primary analysis [47]. b. The cell- and region-level visualization of GT mouse brain WSIs in selected brain slices. For (a) and (b), the 3D/2D reference maps of brain regions are overlaid to aid in the navigation of brain structures. c. The cell- and region-level visualization of generated mouse brain WSIs. d. The visual comparison of exemplar cellular regions ($1024\times 1024$) between Tera-MIND and state-of-the-art (SOTA) patch-based models. From left to right, the image stack displays individual DAPI, PolyT, merged channels, and nuclear masks obtained using Cellpose [29]. e. Quantitative comparison of image generation quality between Tera-MIND and SOTA methods. This includes mean and standard deviation scores for widely-used image quality metrics, such as PSNR and SSIM, alongside domain-specific scores like patch-based nuclear size and cell count. For a more in-depth analysis, we encourage readers to zoom in on the visual results for clearer comparisons.
@@
2.3 Tera-MIND identifies biologically relevant gene-gene interactions within native 3D spatial organizations.
@@
2.3.1 GLUT neuronal system
@@
Glutamate is the primary excitatory neurotransmitter in the brain and plays a pivotal role in a wide range of brain functions [38]. The vesicular glutamate transporters VGLUT1 and VGLUT2, encoded by Slc17a7 and Slc17a6 resp., regulate extracellular glutamate concentrations and synaptic signaling [38, 40]. Moreover, the expression level of VGLUTs impacts the amount of glutamate loaded into vesicles and released, thereby affecting neurotransmission [41]. Changes in VGLUT expression levels have been associated with various neurological pathologies such as schizophrenia, neuropathic pain, and ischemia [24, 34]. In adult mouse brains, spatial expression patterns of VGLUT1 and VGLUT2 are generally complementary but not exclusive  [41]. Specifically, Slc17a7 (VGLUT1) is broadly expressed in the cerebral cortex, cerebellar cortex, hippocampus, and thalamus, etc., while Slc17a6 (VGLUT2) can be found in subcortical regions such as the thalamus and spinal cord. As shown in Fig.3 (a, b), spatial expression patterns of these transporters are accurately mapped in 3D space and align with prior studies [47, 49], which justifies the registration process of brain atlas utilized in this study. Interestingly, the 3D gene-gene attention map learned by Tera-MIND reveals heterogeneous and widespread attention signals for the spatial interaction between Slc17a6 and Slc17a7 (See also Fig. 3 (a) and (d)). Since both transporters collectively contribute to the regulation of synaptic vesicle glutamate content in those overlapped regions, the spatial Slc17a6-Slc17a7 attention signal (Fig. 3 (a, d)) and their positive correlation to the gene expression level (Fig. 3 (c)) likely encapsulate their co-expression patterns and accumulative impact on neurotransmission and regional brain morphology.
2.3.2 DOPA neuronal system

### v3

@@
Tera-MIND: Tera-scale mouse brain simulation via spatial mRNA-guided diffusion
Jiqing Wu Department of Biomedical Engineering, University of Basel, Switzerland Ingrid Berg Department of Pathology and Molecular Pathology, University Hospital, University of Zurich, Switzerland Yawei Li Computer Vision Lab, ETH Zurich, Switzerland Integrated System Laboratory, ETH Zurich, Switzerland Ender Konukoglu Computer Vision Lab, ETH Zurich, Switzerland Viktor H. Koelzer Department of Biomedical Engineering, University of Basel, Switzerland Institute of Medical Genetics and Pathology, University Hospital Basel, Switzerland Correspondence: Viktor.Koelzer@usb.ch
SUMMARY
@@
Holistic 3D modeling of molecularly defined brain structures is crucial for understanding complex brain functions. Using emerging tissue profiling technologies, researchers charted comprehensive atlases of mammalian brain with sub-cellular resolution and spatially resolved transcriptomic data. However, these tera-scale volumetric atlases pose computational challenges for modeling intricate brain structures within the native spatial context. We propose Tera-MIND, a novel generative framework capable of simulating Tera-scale Mouse braINs in 3D using a patch-based and boundary-aware Diffusion model. Taking spatial gene expression as conditional input, we generate virtual mouse brains with comprehensive cellular morphological detail at teravoxel scale. Through the lens of 3D gene-gene self-attention, we identify spatial molecular interactions for key transcriptomic pathways, including glutamatergic and dopaminergic neuronal systems. Lastly, we showcase the translational applicability of Tera-MIND on previously unseen human brain samples. Tera-MIND offers an efficient generative modeling of whole virtual organisms, paving the way for integrative applications in biomedical research.
KEYWORDS
Spatial transcriptomics, mouse and human brain, Generative AI, diffusion model
INTRODUCTION
@@
Recently, a network of researchers 1, 2, 3, 4 has reached a groundbreaking milestone by charting the first complete cellular atlas of the adult mouse brain, the most extensively studied animal model in neuroscience. Using advanced spatial transcriptomics methods 5, 6, 7 capable of profiling molecularly defined brain anatomy and morphology at sub-cellular resolution, the collective studies established foundational resources and datasets, providing an unprecedented framework for systematically exploring the molecular-driven spatial organization of the mammalian brain. Consequently, this landmark achievement opens the door for next-generation investigations into intricate brain functions characterized by exceptional spatial complexity. However, these tera-scale volumetric datasets, including spatial molecular data (e.g., mRNA readouts acting as proxies for gene expression) and paired morphological bioimages, pose substantial computational challenges 8 for in-silico modeling of the mouse brain. Integrating spatial gene expression with 3D brain morphology to uncover functional relationships at whole organ scale remains an open challenge.
@@
Meanwhile, generative artificial intelligence (GenAI) has emerged as a transformative tool for automating visual content creation and manipulation 9, 10. Given a text prompt that describes the desired visual content, GenAI models 11, 12 are able to produce highly realistic images and videos 13, 14 with remarkable efficiency. In biomedical research, the applications of GenAI have gained domain-specific recognition in various areas such as digital histopathology 15, drug screening 16, 17, and molecule design 9. However, the direct application of existing GenAI methods to whole organ simulation has currently remained infeasible due to technological limitations. Developed upon the framework of Generative Adversarial Nets (GAN) 11, the GigaGAN 18 model was trained on 96-128 A100 GPUs to synthesize $4,096\times 4,096$ color images. Using up to 6,144 H100 GPUs, Movie Gen 19, a diffusion-based model 12, achieved video generation results with $1,092\times 1,080\times 256$ spatial-temporal resolution. By contrast, each whole slide image (WSI; $0.108\mu m/\mathrm{pixel}$) of the quality-controlled mouse brain atlas has a spatial resolution of $73,216\times 105,984$. Consequently, stacking 50 slices of these DAPI (4’,6-diamidino-2-phenylindole) and PolyT-stained (total mRNA signals) WSIs results in $0.77\times 10^{12}$ voxels in total. This immense data scale of WSIs and paired gene expression arrays significantly intensifies the computational demands associated with GenAI modeling of a virtual mouse brain.
@@
Addressing memory bottlenecks in high-resolution image generation has previously motivated the development of patch-based approaches 20, 21, 22, 23, 24, which enforce boundary consistency between generated neighboring patches. Similar to many standard GenAI models, these approaches can also leverage the interplay between textual and visual modalities and create desired visual content conditioned on text prompts. While this design choice has shown remarkable success in natural image synthesis and artistic content creation, it is suboptimal for biomedical applications – particularly for the accurate 3D molecular-driven reconstruction of mammalian brains. The inherent complexity of brain morphology controlled by spatial molecular interactions demands a generative framework that can capture molecular-to-morphology spatial associations with precision. In response to these challenges, we propose Tera-MIND, a novel GenAI approach designed for the paired data structure of spatial mRNA arrays (gene expression, prompts) and WSIs (brain morphology, bioimages). Specifically, Tera-MIND employs a newly introduced patch-based and boundary-aware diffusion model, which allows the reconstruction of tera-scale mouse brains with high fidelity. Owing to the patch-wise training and inference paradigm, Tera-MIND models in-silico mammalian brain(s) with computational efficiency. As a result, the whole simulation process can be efficiently executed on a single DGX A100 machine, significantly lowering hardware demands while maintaining scalability for tera-scale data processing. This efficiency underscores the potential wide applicability of Tera-MIND for biomedical applications. Our contributions can be summarized as follows:
@@
  • Similar to the translational aims of wet-lab mouse brain studies, we show that Tera-MIND’s simulation results are applicable to previously unseen human health and glioblastoma brain samples, underscoring its potential in advancing the understanding of complex brain functions and diseases for real-world biomedical applications.
RESULTS
Model overview of Tera-MIND
[…]
@@ RESULTS
Tera-MIND accurately generated tera-scale mouse brain(s) by spatial gene expression.
@@
Consistent with the primary analysis described in 2, which was conducted on a comprehensive brain atlas of a P56 female adult mouse (Sampleid: 638850), we present main simulation results of Tera-MIND based on the same brain atlas. The holistic 3D comparison between the ground truth (GT) and our generation result is illustrated in Fig. 1 (b) and Fig. 2 (a). Given 3D spatial gene expression data as an input, Tera-MIND achieves the faithful and comprehensive reconstruction of 3D brain morphology, as captured by a sequential stack of DAPI and PolyT-stained WSIs. Specifically, the generated mouse brain has the same spatial resolution of 0.77 teravoxels as GT. At the level of individual WSIs (Fig. 2 (b, c)), Tera-MIND exhibits high-quality generation across multiple scales, including cellular, regional, and slice-wise morphological organization. Within distinct brain regions, such as the olfactory areas, isocortex, and cerebellar cortex, Tera-MIND accurately reconstructs complex morphological structures while preserving native cellular distributions. These results highlight the robustness of Tera-MIND for reconstructing 3D brain morphology directly from spatial gene expression profiles.
@@
Complementary to the side-by-side qualitative comparisons, we perform thorough quantitative analyses of the reconstruction quality by Tera-MIND. For a fair and systematic assessment of the biomedical-specific outputs, we report not only commonly used metrics such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM), but also include patch-based morphometric analyses of nuclear size and cell number. These domain-specific metrics are critical for evaluating the biological fidelity of reconstructed brain morphology at the cellular level. Then, we perform quanlitative (Fig. 2 (d)) and quantitative (Fig. 2 (e)) evaluation on Tera-MIND in comparison to state-of-the-art (SOTA) methods CoCoGAN 20, InfinityGAN 21, IST-editing 25, SinFusion 23, Patch-DM 24, etc. Overall, Tera-MIND achieved superior performance in terms of better PSNR and SSIM scores. Furthermore, experimental results reported in Fig. 2 (e) demonstrated marginal morphometric discrepancies between GT cellular patches and those generated by Tera-MIND, which are consistent with the results of stratified patch examples shown in Fig. 2 (e). Using spatial gene expression as the input prompt, these analyses confirm that Tera-MIND reliably reconstructs brain morphology across multiple scales, from cellular and regional structures to whole-brain organization. This consistent performance across scales highlights the efficacy of our approach for accurately capturing the complex spatial patterns of 3D brain morphology.
@@
Tera-MIND identified biologically relevant gene-gene interactions in native 3D space.
@@
GLUT neuronal system
@@
Glutamate is the primary excitatory neurotransmitter in the brain and plays a pivotal role in a wide range of brain functions 26. The vesicular glutamate transporters VGLUT1 and VGLUT2, encoded by Slc17a7 and Slc17a6 resp., regulate extracellular glutamate concentrations and synaptic signaling 26, 27. Moreover, the expression level of VGLUTs impacts the amount of glutamate loaded into vesicles and released, thereby affecting neurotransmission 28. Changes in VGLUT expression levels are associated with various neurological pathologies such as schizophrenia, neuropathic pain, and ischemia 29, 30. In adult mouse brains, spatial expression patterns of VGLUT1 and VGLUT2 are generally complementary but not exclusive  28. Specifically, Slc17a7 (VGLUT1) is broadly expressed in the cerebral cortex, cerebellar cortex, hippocampus, and thalamus, etc., while Slc17a6 (VGLUT2) can be found in subcortical regions such as the thalamus and spinal cord.
As shown in Fig. 3 (a, left) and (b), spatial expression patterns of these transporters are accurately mapped in 3D space and align with prior studies 1, 2, which justifies the registration process of brain atlas utilized in this study. Interestingly, the 3D gene-gene attention map (Fig. 3 (a, middle)) learned by Tera-MIND reveals widespread and heterogeneous attention signals (Fig. 3 (a, right)) for the spatial interaction between Slc17a6 and Slc17a7. Such heterogeneous attention levels can also be observed by the marginal distribution reported in Fig. 3 (c), where the linear regression analysis identifies a strong correlation between expression and attention levels. Since both transporters collectively contribute to the regulation of synaptic vesicle glutamate content in those overlapped regions, these heterogeneous spatial attention signals (Fig. 3 (c, d)) likely encapsulate their accumulative impact on regional brain morphology and neurotransmission.
DOPA neuronal system